# SALMONN: Towards Generic Hearing Abilities for Large Language Models

### v1

@@ Abstract
Hearing is arguably an essential ability of artificial intelligence (AI) agents in the physical world, which refers to the perception and understanding of general auditory information consisting of at least three types of sounds: speech, audio events, and music. In this paper, we propose SALMONN, a speech audio language music open neural network, built by integrating a pre-trained text-based large language model (LLM) with speech and audio encoders into a single multimodal model. SALMONN enables the LLM to directly process and understand general audio inputs and achieve competitive performances on a number of speech and audio tasks used in training, such as automatic speech recognition and translation, auditory-information-based question answering, emotion recognition, speaker verification, and music and audio captioning etc. SALMONN also has a diverse set of emergent abilities unseen in the training, which includes but is not limited to speech translation to untrained languages, speech-based slot filling, spoken-query-based question answering, audio-based storytelling, and speech audio co-reasoning etc. The presence of the cross-modal emergent abilities is studied, and a novel few-shot activation tuning approach is proposed to activate such abilities of SALMONN. To our knowledge, SALMONN is the first model of its type and can be regarded as a step towards AI with generic hearing abilities. An interactive demo of SALMONN is available at https://github.com/bytedance/SALMONN, and the training code and model checkpoints will be released upon acceptance.
@@ 3.2 Training Method
Task Over-fitting: Although SALMONN built with only the first two stages of training can produce competitive results on the tasks trained in instruction tuning, it exhibits limited or almost no ability to perform untrained cross-modal tasks, especially the tasks that require cross-modal co-reasoning abilities. In particular, the model sometimes violates the instruction prompts and generates irrelevant responses as if it received an instruction related to a task commonly seen in training (e.g. speech recognition). We refer to this phenomenon as task over-fitting. A theoretical analysis of the issue is presented here and detailed experimental verification is provided later in Section 5.4.
We attribute task over-fitting to two reasons. First, compared to the text-only data used in LLM training, only simpler instruction prompts are used in our cross-modal instruction tuning (Wei et al., 2022a), and the resulting responses are not as complex and diverse. Meanwhile, some tasks included in instruction tuning, in particular speech recognition and audio captioning, have more deterministic outputs than the other tasks, such as speech and audio question answering. These two reasons combined to cause the intrinsic conditional language model (LM) to bias to an ill-formed distribution with poor generalisation ability, which prohibits SALMONN from performing untrained cross-modal tasks. More specifically, denote $\{\mathbf{I}_{i}\}_{i=1}^{I}$ as the set of all the training prompts and for each prompt, there is a set of paired training samples $\{\mathbf{X}_{ij},\mathbf{Y}_{ij}\}_{j=1}^{J_{i}}$, where $\mathbf{X}_{ij}$ and $\mathbf{Y}_{ij}$ are the audio input and text supervision of prompt $\mathbf{I}_{i}$. Therefore, our training objective is
$\hat{\mathbf{\Lambda}}=\arg\max_{\mathbf{\Lambda}}\sum\nolimits_{i=1}^{I}\sum\nolimits_{j=1}^{J_{i}}P_{\mathbf{\Lambda}}(\mathbf{Y}_{ij}|\mathbf{X}_{ij},\mathbf{I}_{i}),$ (3)
where $\mathbf{\Lambda}$ refers to a set of model parameters. At test time, the response text sequence $\hat{\mathbf{Y}}$ of a test input $\mathbf{X}$ given a new instruction prompt $\mathbf{I}$ can be generated according to
$\hat{\mathbf{Y}}=\arg\max\nolimits_{\mathbf{Y}}P_{\mathbf{\Lambda}}(\mathbf{Y}|\mathbf{X},\mathbf{I}).$ (4)
Using the Bayes’ Rule, there is
$P_{\mathbf{\Lambda}}(\mathbf{Y}|\mathbf{X},\mathbf{I})=\frac{P_{\mathbf{\Lambda}}(\mathbf{I}|\mathbf{Y},\mathbf{X})P_{\mathbf{\Lambda}}(\mathbf{Y}|\mathbf{X})}{P_{\mathbf{\Lambda}}(\mathbf{I}|\mathbf{X})}.$ (5)
From Eqn. (3), it is clear that training may fit better with some prompts $\mathbf{I}^{{}^{\prime}}$ that correspond to more training samples or lead to higher $P_{\mathbf{\Lambda}}(\mathbf{Y}|\mathbf{X},\mathbf{I}^{{}^{\prime}})$, such as the prompts used for speech recognition and audio captioning. Furthermore, since only limited text responses are seen in SALMONN training, the intrinsic conditional LM $P_{\mathbf{\Lambda}}(\mathbf{Y}|\mathbf{X})$ is biased towards simple and short responses. From Eqn. (5), this causes $\mathbf{I}^{\prime\prime}$, zero-shot instructions with more diverse responses, to have small $P_{\mathbf{\Lambda}}(\mathbf{Y}|\mathbf{X},\mathbf{I}^{\prime\prime})$.
@@ 4.1 Model Specifications
SALMONN uses the encoder part of Whisper-Large-v2111https://huggingface.co/openai/whisper-large-v2 (Radford et al., 2023) model as the speech encoder, the fine-tuned BEATs [^2] Fine-tuned BEATs_iter3+ (AS2M) (cpt2) (Chen et al., 2023c) encoder as the audio encoder, and a Vicuna LLM with 13 billion parameters [^3] https://huggingface.co/lmsys/vicuna-13b-v1.1 (Chiang et al., 2023) as the backbone LLM. For the window-level Q-Former, we use $N=1$ resulting in only one trainable query, and use $L=17$ which is approximately 0.33 seconds per window. This leads to 88 textual tokens output by Q-Former for a 30-second audio. Regarding the hyper-parameters of LoRA (Hu et al., 2022), we set the rank to 8 and the scaling factor to 4.0. Only the parameters of Q-Former and LoRA are updated in training, which resulted in $\sim$ 33 million (M) parameters that is $\sim$0.24of the entire SALMONN model.
@@ 4.2 Data specifications
The second instruction tuning stage involves multiple tasks, including automatic speech recognition (ASR), automatic speech translation (AST), automatic audio captioning (AAC), phone recognition (PR), emotion recognition (ER), music captioning (MC), overlapped speech recognition (OSR), speaker verification (SV), gender recognition (GR) and speech question answering (SQA), audio question answering (AQA) and music question answering (MQA). In the SQA, AQA and MQA tasks, the questions are generated based on the text caption labels using GPT-3.5444https://chat.openai.com, and the model needs to provide answers based on the general audio input and the text prompt with a question. The data used in this stage is listed in Table 1, where “En2Zh” refers to AST from English to Chinese.
@@ 4.3 Task Specifications
Task level 3 has the most difficult tasks including audio-based storytelling (Story) and speech audio co-reasoning (SAC). Audio-based storytelling is to write a meaningful story based on the auditory information from the general audio inputs. SAC requires the model to understand a spoken question embedded in the input audio clip, find evidence from the background audio events or music, and reason from it to answer the question. Both level 3 tasks are new tasks that first proposed in this paper, to the best of our knowledge, which require SALMONN to perceive speech, audio, and music, and to understand and reason based on the auditory information in a fully end-to-end way.
@@
Table 2: Test sets, metrics, and sources of the reference values used in the three levels of tasks. The speech data used in SQQA and KE are synthesised using a commercial text-to-speech product. For reference values: “Whisper” refers to using Whisper ASR, “Whisper + Vicuna” refers to feeding Whisper output text into Vicuna, and the rest are from literature.
@@
En2Zh CoVoST2-En2Zh BLEU4 Whisper + Vicuna
AAC AudioCaps METEOR |||| SPIDEr SOTA (Mei et al., 2023)
@@
MC MusicCaps BLEU4, RougeL SOTA (Doh et al., 2023)
@@ 5.1 Full results on all 15 tasks
w/o Activation (2.1, 4.9, 9.1) 34.4 25.6 |||| 47.6 17.6 0.63 3.5, 22.1 20.7 0.93
w/ Activation (2.1, 4.9, 10.0) 33.1 24.0 |||| 40.3 17.6 0.69 5.5, 21.8 23.0 0.94
Reference Value (2.2, 5.1, 9.2) 14.6 25.0 |||| 48.5 3.1 0.81 6.1, 21.5 7.6 -
@@ 5.2 Discounting Lora scaling factor
This section explores the influence of the use of test-time discounting of the LoRA scaling factor for alleviating the task over-fitting issue. As shown in Fig. 3, when LoRA scaling factor decreases to around 2.0, the model suddenly emerges with cross-modal reasoning abilities, which, together with the drops in %PER, proves the existence of the intrinsic conditional LM embedded in LoRA.
@@
5.3 Analysis of task over-fitting
Model Checkpoint ASR AAC PR Story SAC
@@ 5.4 Activation tuning using different tasks and data
The results are shown in Table 5. From the results, neither ASR (Long), training ASR with long labels, nor Story (Text based), training LoRA based on only text prompt input, can activate the model. Training with stories or QA with long answers, Story or QA (Long), can. This possibly lies in the fact that the ASR task emphasises improving high-quality cross-modal alignments that make the distribution even more biased towards the intrinsic conditional LM. As for text-based fine-tuning on stories, this actually affect $P(\mathbf{Y}|\mathbf{T}_{\text{x}},\mathbf{I})$ instead of $P(\mathbf{Y}|\mathbf{X},\mathbf{I})$, where $\mathbf{T}_{\text{x}}$ is the reference text of the speech. Therefore, text-based training cannot alleviate task over-fitting. QA data with long answers can activate the LLM, but the performance of the activated model is worse than that activated using the Story task, which especially results in a higher repeat rate of the response generation. That is possibly due to the fact that the paired answer to the QA task is less diverse than the Story task, which leads to less effective activation of the cross-modal emergent abilities.
@@
w/o Activation 2.1 17.6 0.19 (0.29) 7.77 (0.00) 0.02 (0.04) 0.2%
Story 2.1 17.6 0.41 (0.98) 82.57 (1.00) 0.50 (0.73) 0.1%
QA (Long) 2.1 17.6 0.40 (0.93) 59.82 (1.00) 0.34 (0.71) 4.6%
ASR (Long) 2.2 17.7 0.22 (0.28) 7.87 (0.00) 0.12 (0.03) 0.1%
Story (Text based) 2.1 17.6 0.23 (0.32) 8.45 (0.03) 0.11 (0.03) 0.1%
Story (LoRA only) 2.1 17.6 0.44 (0.96) 82.29 (1.00) 0.34 (0.65) 0.2%
@@ 7 Reproducibility Statement
To make the experiments and models reproducible, the training data and the benchmark details are provided in Section 4.2 and Section 4.3. The URL of an interactive demo page is provided in the abstract for a convenient try-out of the model. The complete training and inference code together with model checkpoints will be released upon acceptance.
@@ Appendix A Examples of SALMONN
SALMONN is capable of a large range of auditory tasks. Here we list some examples in Fig. 4-20.
@@
Appendix B Changes in Perplexity during Training
Changes in perplexity during the proposed three-stage training on six tasks including ASR, AAC, AST (En2Zh), OSR, Story, SAC are visualized in Fig. 21.

### v2

@@ Abstract
Hearing is arguably an essential ability of artificial intelligence (AI) agents in the physical world, which refers to the perception and understanding of general auditory information consisting of at least three types of sounds: speech, audio events, and music. In this paper, we propose SALMONN, a speech audio language music open neural network, built by integrating a pre-trained text-based large language model (LLM) with speech and audio encoders into a single multimodal model. SALMONN enables the LLM to directly process and understand general audio inputs and achieve competitive performances on a number of speech and audio tasks used in training, such as automatic speech recognition and translation, auditory-information-based question answering, emotion recognition, speaker verification, and music and audio captioning etc. SALMONN also has a diverse set of emergent abilities unseen in the training, which includes but is not limited to speech translation to untrained languages, speech-based slot filling, spoken-query-based question answering, audio-based storytelling, and speech audio co-reasoning etc. The presence of cross-modal emergent abilities is studied, and a novel few-shot activation tuning approach is proposed to activate such abilities. To our knowledge, SALMONN is the first model of its type and can be regarded as a step towards AI with generic hearing abilities. The source code, model checkpoints and data are available at https://github.com/bytedance/SALMONN.
@@ 3.2 Training Method
Task Over-fitting: Although SALMONN built with only the first two stages of training can produce competitive results on the tasks trained in instruction tuning, it exhibits limited or almost no ability to perform untrained cross-modal tasks, especially the tasks that require cross-modal co-reasoning abilities. In particular, the model sometimes violates the instruction prompts and generates irrelevant responses as if it received an instruction related to a task commonly seen in training (e.g. speech recognition). We refer to this phenomenon as task over-fitting. A theoretical analysis of the issue is presented here and detailed experimental verification is provided in Sections 5.3 and 5.4.
We attribute task over-fitting to two reasons. First, compared to the text-only data used in LLM training, only simpler instruction prompts are used in our cross-modal instruction tuning (Wei et al., 2022a) and the resulting responses are not as complex and diverse. Meanwhile, some tasks included in instruction tuning, in particular speech recognition and audio captioning, have more deterministic outputs than the other tasks, such as speech and audio question answering. These two reasons combined to cause the intrinsic conditional language model (LM) to bias to an ill-formed distribution with poor generalisation ability, which prohibits SALMONN from performing untrained cross-modal tasks. More specifically, at test time the response text sequence $\hat{\mathbf{Y}}$ of a test input $\mathbf{X}$ given a new instruction prompt $\mathbf{I}$ can be generated according to $\hat{\mathbf{Y}}=\arg\max\nolimits_{\mathbf{Y}}P_{\mathbf{\Lambda}}(\mathbf{Y}|\mathbf{X},\mathbf{I})$, which is also the objective to maximise in training. Using the Bayes’ Rule, there is
$P_{\mathbf{\Lambda}}(\mathbf{Y}|\mathbf{X},\mathbf{I})=P_{\mathbf{\Lambda}}(\mathbf{Y}|\mathbf{X})\cdot P_{\mathbf{\Lambda}}(\mathbf{I}|\mathbf{Y},\mathbf{X})/{P_{\mathbf{\Lambda}}(\mathbf{I}|\mathbf{X})}.$ (3)
Since only limited text responses are seen in SALMONN training, the intrinsic conditional LM $P_{\mathbf{\Lambda}}(\mathbf{Y}|\mathbf{X})$ is biased towards the $\mathbf{Y}$ sequences that are strongly-aligned with $\mathbf{X}$, such as the transcriptions of automatic speech recognition (ASR) and automatic audio captioning (AAC) tasks, especially simple and short transcriptions. From Eqn. (3), this causes $\mathbf{I}^{\prime}$, zero-shot instructions with more diverse responses, to have small $P_{\mathbf{\Lambda}}(\mathbf{Y}|\mathbf{X},\mathbf{I}^{\prime})$.
@@ 4.1 Model Specifications
SALMONN uses the encoder part of Whisper-Large-v2 (Radford et al., 2023) model as the speech encoder, the fine-tuned BEATs (Chen et al., 2023c) encoder as the audio encoder, and a Vicuna LLM with 13 billion parameters (Chiang et al., 2023) as the backbone LLM. For the window-level Q-Former, we use $N=1$ resulting in only one trainable query, and use $L=17$ which is approximately 0.33 seconds per window. This leads to 88 textual tokens output by Q-Former for a 30-second audio. Regarding the hyper-parameters of LoRA (Hu et al., 2022), we set the rank to 8 and the scaling factor to 4.0. Only the parameters of Q-Former and LoRA are updated in training, which resulted in $\sim$ 33 million (M) parameters that is $\sim$0.24of the entire SALMONN model.
@@ 4.2 Data specifications
The second instruction tuning stage involves multiple tasks, including ASR, automatic speech translation (AST), AAC, phone recognition (PR), emotion recognition (ER), music captioning (MC), overlapped speech recognition (OSR), speaker verification (SV), gender recognition (GR) and speech question answering (SQA), audio question answering (AQA) and music question answering (MQA). In the SQA, AQA and MQA tasks, the questions are generated based on the text caption labels using ChatGPT, and the model needs to provide answers based on the general audio input and the text prompt with a question. The data used in this stage is listed in Table 1, where “En2Zh” refers to AST from English to Chinese.
@@ 4.3 Task Specifications
Task level 3 has the most difficult tasks including audio-based storytelling (Story) and speech audio co-reasoning (SAC). Audio-based storytelling is to write a meaningful story based on the auditory information from the general audio inputs. SAC requires the model to understand a spoken question embedded in the input audio clip, find evidence from the background audio events or music, and reason from it to answer the question. Both level 3 tasks are new tasks that are first proposed in this paper, to the best of our knowledge, which requires SALMONN to perceive speech, audio, and music, and to understand and reason based on the auditory information in a fully end-to-end way.
@@
Table 2: Test sets, metrics, and sources of the reference values used in the three levels of tasks. The speech data used in SQQA and KE are synthesised using a commercial text-to-speech product. For reference values: “Whisper” refers to using Whisper-Large-v2 for ASR, “Whisper + Vicuna” refers to feeding the Whisper-Large-v2 ASR output transcriptions into Vicuna, and the rest are the state-of-the-art results to the best of our knowledge.
@@
En2Zh CoVoST2-En2Zh BLEU4 (Wang et al., 2021)
AAC AudioCaps METEOR |||| SPIDEr (Mei et al., 2023)
@@
MC MusicCaps BLEU4, RougeL (Doh et al., 2023)
@@ 5.1 Full results on all 15 tasks
More detailed analysis about the results are discussed in Appendix B due to page limit.
@@
w/o Activation (2.1, 4.9, 9.1) 34.4 25.6 |||| 47.6 4.2 0.63 3.5, 22.1 20.7 0.93
w/ Activation (2.1, 4.9, 10.0) 33.1 24.0 |||| 40.3 4.2 0.69 5.5, 21.8 23.0 0.94
Reference Value (2.2, 5.1, 9.2) 38.9 25.0 |||| 48.5 3.1 0.81 6.1, 21.5 7.6 -
@@ 5.2 Discounting Lora scaling factor
This section explores the influence of the use of test-time discounting of the LoRA scaling factor for alleviating the task over-fitting issue without activation tuning. As shown in Fig. 3, when the LoRA scaling factor decreases to around 2.0, i.e. to half of its original value, the model suddenly emerges with cross-modal reasoning abilities, which, together with the drops in %PER, proves the existence of the intrinsic conditional LM embedded in LoRA.
@@
5.3 Analysis of task over-fitting and activation tuning
To verify the influences of the intrinsic conditional LM, we calculate the perplexities (PPLs) of $P_{\mathbf{\Lambda}}(\mathbf{Y}|\mathbf{X},\mathbf{I})$ and $P_{\mathbf{\Lambda}}(\mathbf{Y}|\mathbf{X})$ of each step of the activation tuning stage. In detail, for a given audio $\mathbf{X}$, the probability of the $\mathbf{Y}$ sequence corresponding to the probed task is calculated using teacher forcing based on the text instruction prompt $\mathbf{I}$ of a certain task or without using any $\mathbf{I}$.
As shown in Fig. 4, PPLs were compared between the ground-truth responses of the task of Story/SAC (generated by discounting the LoRA scaling factor) and the responses of the task that the model performed incorrectly due to task over-fitting (AAC in this example). In sub-figures (a) and (b), no text instruction prompt was used when computing $P_{\mathbf{\Lambda}}(\mathbf{Y}|\mathbf{X})$, showing that without activation tuning the PPL of the $\mathbf{Y}$ of AAC was obviously lower than that of the $\mathbf{Y}$ of Story/SAC. During activation tuning, the PPL gaps between these tasks were mitigated, meaning the bias to the dominant AAC task was considerably reduced. In sub-figures (c) and (d), we probed $P_{\mathbf{\Lambda}}(\mathbf{Y}|\mathbf{X},\mathbf{I})$, the PPLs with the instructions of Story and SAC respectively. It is revealed that before activation tuning, the PPL of AAC is lower than that of Story/SAC even if the text instruction prompt is to perform Story/SAC, which explains the failure of the model to follow the instruction. During activation tuning, the PPL values of the $\mathbf{Y}$ of Story/SAC gradually become lower than those of AAC, and the model can eventually perform the instructed task.
[…]
Figure 4: Changes in PPL during the activation tuning stage. (a) shows $-\ln P_{\mathbf{\Lambda}}(\mathbf{Y}|\mathbf{X})$ for AAC and Story, (b) shows $-\ln P_{\mathbf{\Lambda}}(\mathbf{Y}|\mathbf{X})$ for AAC and SAC, (c) shows $-\ln P_{\mathbf{\Lambda}}(\mathbf{Y}|\mathbf{X},\mathbf{I}=\text{Story})$, and (d) shows $-\ln P_{\mathbf{\Lambda}}(\mathbf{Y}|\mathbf{X},\mathbf{I}=\text{SAC})$.
@@ 5.4 Activation tuning using different tasks and data
The results are shown in Table 4. From the results, neither ASR (Long), training ASR with long labels, nor Story (Text based), training LoRA based on only text prompt input, can activate the model. Training with stories or QA with long answers, Story or QA (Long), can. This lies in the fact that the ASR task emphasises improving high-quality cross-modal alignments that make the distribution even more biased towards the intrinsic conditional LM. As for text-based fine-tuning on stories, this actually affect $P(\mathbf{Y}|\mathbf{T}_{\text{x}},\mathbf{I})$ instead of $P(\mathbf{Y}|\mathbf{X},\mathbf{I})$, where $\mathbf{T}_{\text{x}}$ is the reference text of the speech. Therefore, text-based training cannot alleviate task over-fitting. QA data with long answers can activate the LLM, but the performance of the activated model is worse than that activated using the Story task, which especially results in a higher repeat rate of the response generation. That is possibly due to the fact that the paired answer to the QA task is less diverse than the Story task, which leads to less effective regularisation of the intrinsic conditional LM $P_{\mathbf{\Lambda}}(\mathbf{Y}|\mathbf{X})$.
@@
w/o Activation 2.1 4.2 0.19 (0.29) 7.77 (0.00) 0.02 (0.04) 0.2%
Story 2.1 4.2 0.41 (0.98) 82.57 (1.00) 0.50 (0.73) 0.1%
QA (Long) 2.1 4.3 0.40 (0.93) 59.82 (1.00) 0.34 (0.71) 4.6%
ASR (Long) 2.2 4.2 0.22 (0.28) 7.87 (0.00) 0.12 (0.03) 0.1%
Story (Text based) 2.1 4.2 0.23 (0.32) 8.45 (0.03) 0.11 (0.03) 0.1%
Story (LoRA only) 2.1 4.2 0.44 (0.96) 82.29 (1.00) 0.34 (0.65) 0.2%
@@ 7 Reproducibility Statement
To make the experiments and models reproducible, the training data and the benchmark details are provided in Section 4.2 and Section 4.3. The source code, model checkpoints, and data are released on the SALMONN project GitHub page.
@@ Appendix A Examples of SALMONN
SALMONN is capable of a large range of auditory tasks. Here we list some examples in Fig. 5-21.
@@
Appendix B Performance Analysis
As shown in Table 3(b), with activation tuning [^2] Added at the requests of anonymous reviewers. We thank the reviewers for the suggestion., SALMONN
  1. 1.
    performs close to the state-of-the-art results on the trained tasks including ASR, En2Zh, AAC and MC;
  2. 2.
    learns to handle PR and OSR tasks which are difficult for the cascaded approach of “Vicuna + Whisper”;
  3. 3.
    generalises well on a wide range of speech-grounded NLP tasks, especially such as En2De and En2Ja where SALMONN performs better than the ”Whisper+Vicuna” cascaded system, which indicates the superior of the audio-grounding system to cascade system on avoiding error propagation and the loss of non-linguistic information (e.g. prosody);
  4. 4.
    tackles tasks such as audio-based storytelling and SAC which existing models can not handle to our best knowledge.
The underlying reason for using the cascaded Whisper + Vicuna system for reference values of the level 2 tasks lies in the fact that all level 2 tasks are zero-shot and there is no other audio-grounding system apart from SALMONN that can perform such tasks as zero-shot. An advantage of SALMONN over the cascaded system is that SALMONN can leverage both linguistic and paralinguistic information required for spoken language understanding.
Despite such advantages, SALMONN has performance limitations on some tasks.
  1. 1.
    First, PR is achieved by extending the LLM to consider phonemes as a new writing system. Since recognising phonemes requires finer-grained modelling of pronunciation information than recognising the word pieces used by the original Whisper ASR, it is not easy for the SALMONN model built upon an existing Whisper speech encoder to perform as well as a specialised model on the PR task.
  2. 2.
    Similarly, SALMONN has a relatively high WER on OSR since the Whisper ASR was not able to perform OSR.
  3. 3.
    The success of SQQA mainly relies on the understanding of the spoken questions (e.g. “What is the highest mountain in the world”) and answering the questions based on the commonsense knowledge stored in the text-based LLM. The drop in SQQA performance indicates that the use of LoRA cross-modal adaptation may cause the LLM to “forget” some text-based commonsense knowledge.
Appendix C Statistics of task over-fitting
Model Checkpoint ASR AAC PR Story SAC
Appendix D Changes in Perplexity during Training
Changes in perplexity during the proposed three-stage training on six tasks including ASR, AAC, AST (En2Zh), OSR, Story, and SAC are visualized in Fig. 22.